what is the difference between a trig equation that is an identity and a trig equations tha tis not an identity? Provide an example to clarify.
Answer by stanbon(48544)

A trig identity is true for all values of x,
e.g. sin x = 1/ csc x
A trig equation that is not an identity is true
only for certain values of x
e.g. sin x = 1/2 for x = pi/6 +- 2kpi or x = 5pi/6 +- 2kpi, where k=0,1,2,... That's a lot of values but it is not "all possible values".
Cheers,
Stan H.